OpenStudy (anonymous):

find a cartesian cuestion r=2cos(theta) + 3sin(theta)

OpenStudy (anonymous):

x/r = cos(theta) y/r = sin(theta) r^2 = x^2 +y^2 have you made those substitutions?

OpenStudy (anonymous):

help me you please.... with this

OpenStudy (anonymous):

make those substitutions... tell me what you get...

OpenStudy (anonymous):

r= 2x/r + 3 y/r ...........> r2=2x+2y .. ...............> x2 + y2= 2x + 3y ------> x2-2x-3y+y2= 0... is it correct¿?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can clean it up a bit by completing the square on both the x and y terms.. eg.: (x^2 -2x) = (x^2 -2x +1) -1 = (x-1)^2 -1

OpenStudy (anonymous):

what is the best¿?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

ok ok thanks .... can you help with other¿?

OpenStudy (anonymous):

xy=8

OpenStudy (anonymous):

just sub.s in x=rcos(theta), y=rsin(theta)

OpenStudy (anonymous):

rcostheta + rsintheta = 8 ¿?